OpenStudy (anonymous):

My teacher gave me this problem: 3 + 2 = 1 x-3 x+3 The answer he gave us is: x= -1 +/- (square root) 217 6 he wants us to show a check... please explain to me how to check this answer >(

OpenStudy (anonymous):

 drunk frog for encouragement >;3

OpenStudy (anonymous):

oh my gosh :D Thanks, I will need It with the insane teacher I have.

OpenStudy (anonymous):

Figured you did :3

OpenStudy (anonymous):

My brain hurts...

OpenStudy (jdoe0001):

\(\bf 3+\cfrac{2}{x-3}=\cfrac{1}{x+3}\quad ?\)

OpenStudy (anonymous):

not sure what all of that means....

OpenStudy (jdoe0001):

hmmm I see you're using ... Internet Explorer.... got any other browser handy? try to check the same page in another browser

OpenStudy (anonymous):

oh... that makes more sense... yes that is the correct problem

OpenStudy (jdoe0001):

ok,let us move the right-hand side over and you'd get \(\bf 3+\cfrac{2}{x-3}=\cfrac{1}{x+3}\implies 3+\cfrac{2}{x-3}-\cfrac{1}{x+3}=0\\ \quad \\ \textit{let us get the LCD, which will be} (x-3)(x+3)\qquad thus\\ \quad \\ \cfrac{3(x-3)(x+3)+2(x+3)-1(x-3)}{(x-3)(x+3)}=0\)

OpenStudy (jdoe0001):

then multiplying by (x-3)(x+3) both sides we get \(\bf \cfrac{3(x-3)(x+3)+2(x+3)-1(x-3)}{(x-3)(x+3)}=0\\ \quad \\ 3(x-3)(x+3)+2(x+3)-1(x-3)=0\\ \quad \\ \implies 3x^2-27+2x+6-x-3=0\implies 3x^2+x-24=0\) as you can see is a quadratic, so you can either factor it, or use the quadratic formula

OpenStudy (jdoe0001):

hmmm I have one tiny error... it should be -x +3

OpenStudy (anonymous):

i understand how to find the answer. i just need to know how to show a check

OpenStudy (jdoe0001):

\(\bf 3(x-3)(x+3)+2(x+3)-1(x-3)=0\\ \quad \\ \implies 3x^2-27+2x+6-x+3=0\implies 3x^2+x-18=0\)

OpenStudy (jdoe0001):

ohhh just to check is correct well that's simple, just take your values for "x" and plug them in the original equation the left-side and the right-side must equate, thus is a EQUATion

OpenStudy (anonymous):

I know that i have to substitute x for the answer that i found, but i have no clue where to start